# Online Gaussian Process Distance Field with VDB Structure

### v2

@@ IV-D Global Gaussian Process Signed Distance Field
After the fusion process, the leaf nodes in the global VDB grid are marked as active nodes. We perform the marching cubes for the voxels of active nodes and for computing the dense reconstruction with surface properties. The zero crossing points in current active nodes $\mathcal{P}_{\mathcal{S}\{i\}}$ then merge into $\mathcal{P}_{\mathcal{S}\{0,...,i\}}$. Zero crossing points in each node are used to train separate GPs. The combination of all GPs forms the global Gaussian Process Signed Distance Field (G-GPDF).
@@
where $\lambda$ is a large positive value to control the evenness of the smooth minimum. We use $100$ for $\lambda$ in this paper. ${\hat{d}}_{q}(\mathbf{x}_{*})$ is the distance inference at querying location $\mathbf{x}_{*}$ given the located GP node with the index of $q$. For the gradient, we directly average the inferred gradients from the $Q$ clustered GP nodes.
Figure 3: Quantitative comparisons of a) and b) the efficiency performance, c) and d) reconstruction accuracy in Chamfer distance, e) and f) the distance field accuracy on the cow and lady RGB-D dataset and the newer college LiDAR dataset.
@@ V Evaluation
To evaluate the proposed VDB-GPDF framework, we quantitatively compare the mapping performance for a) efficiency, b) reconstruction accuracy and c) distance accuracy, and qualitatively demonstrate the mapping performance for a room size with RGB-D cameras and a LiDAR large-scale dataset. Our framework is implemented in C++ based on ROS1. All experiments were run on 12th Gen Intel® Core™ i5-1245U with 12 cores.
The first dataset is the Cow and Lady [^1] https://projects.asl.ethz.ch/datasets/doku.php?id=iros2017, which includes fibreglass models of a large cow and a lady standing side by side in a room. The dataset consists of RGB-D point clouds collected using a Kinect 1 camera, with sensor trajectory captured using a Vicon motion system. We use frames 220-820 of the dataset, which covers major information of the scene. For the second dataset, we choose the Newer College dataset [25] collected by a handheld sensor platform with a 128-beam ouster LiDAR. The estimated poses are calculated via the alignment of each frame to the ground truth map. The size of the main quad is roughly $56\times 40\times 20$ m, with approximately 2000 frames. Both Cow and Lady and Newer College have the ground truth map that allows us to perform a proper quantitative evaluation for both reconstruction and distance field accuracy. For a fair comparison, we process every frame from both datasets in each framework. Our strategy for length scale selection depends on the number of voxels in each leaf node, which in our case is fixed to $2^{3}\times 2^{3}\times 2^{3}$ voxels. A length scale that correlates two to four voxels is a good trade-off between smoothing and interpolation for our kernel function.
@@ V-B Efficiency
We first evaluate the VDB-GPDF efficiency against the state-of-the-art mapping frameworks VDB-Fusion and VDB-EDT, both developed based on OpenVDB data structure. Fig. LABEL:cow:time and LABEL:newer:time show the average computational time per frame for the room size RGB-D Cow and Lady dataset and the outdoor LiDAR Newer College dataset, respectively. Note that the Newer College quad has pedestrians walking by as dynamic objects. Both frameworks have free-space carving methods enabled to update the moving objects in the scene. For the sake of simplicity, we compute the average time for all the integration processes, including fusion and ESDF. As for our framework, we include L-GPDF, incremental fusion and G-GPDF. Both figures show that our VDB-GPDF is more efficient than other frameworks with varying resolutions due to our voxelisation method and fewer testing points in the frustum area. VDBFusion consumes more time in the RGB-D dataset than others due to enabling the free-space carving process. For the LiDAR data, VDB-EDT requires more time to update the EDF every frame.
@@ V-C Reconstruction Accuracy
Figure 5: Qualitative reconstruction comparison of a) VDBFusion and b) our proposed method shows that VDB-GPDF produces a more complete and natural reconstruction, especially when the sensor is in parallel with the observed surface. Our mesh covers on the top of the mattress, and the colour is properly fused for visualisation as the scene photo colour in c).
VDBFusion and VDB-GPDF both produce similar meshes of the scene shown in Fig. LABEL:cow_mesh_vdbfusion and Fig. LABEL:cow_mesh_ours, respectively. VDB-GPDF produces a more complete and natural reconstruction, especially when the sensor is parallel to the observed surface. It can be seen that there is a gap at the top of the mattresses on the left-hand side in the VDBFusion result, whereas VDB-GPDF does not have the same gap. In addition, our colour is nicely fused with a sample of the RGB camera data in c) for qualitative comparison. Note that, for example, around the wall and floor, our method shows the colour with high fidelity, while VDBFusion colour is more approximate.
We use the Chamfer distance metrics [26] against VDB-Fusion and VDB-EDT to compare the maps on both datasets quantitatively. As demonstrated in Fig. LABEL:cow:chamfer and LABEL:newer:chamfer, our map quality outperforms VDB-EDT in different resolutions. Due to our voxelisation method being applied, we have comparable results to VDBFusion in reconstruction accuracy. To show the performance in a large-scale dataset, we use the main quad of the Newer College, which is roughly $56\times 40\times 20$ m, with approximately 2000 frames in the dataset. In Fig. 1, we qualitatively show our mesh coloured by fused LiDAR intensity and the distance field output. Fig. LABEL:teaser1 shows the incrementally built dense reconstruction. We zoomed in Fig. LABEL:teaser2 to show the textures on the ceiling inside the corridor and Fig. LABEL:teaser3 the stairs and windows at the corner of the quad.
Figure 6: Comparison of ground truth and estimated EDF with a 2D horizontal slice 0.9 m above the ground for the Cow and Lady dataset. The resolution is 0.05m. e) is accurate and closely resembles the ground truth in a). b). c) and d) do not produce enough smoothness in the field and only offer values in the observed area. Our VDB-GPDF with G-GPDF in j) produces lower and smoother values of distance errors than VDB-EDT in g) and Voxblox in h) and our method without the G-GPDF in i).
Figure 7: Comparison of ground truth and estimated EDF with a 2D horizontal slice 0.9 m above the ground for the newer college. d) is accurate and closely resembles the ground truth in a). b) fails to update in the areas where moving objects are present. c) does not produce enough smoothness in the field. Similarly, Our VDB-GPDF in h) produces lower and smoother values of distance errors than VDB-EDT in f) and Voxblox in g) with the LiDAR dataset.
@@ V-D Distance Accuracy
We evaluate the proposed GP distance field with VDB structure against Voxblox and VDB-EDT since the VDBFusion framework has no EDF. We use the ground truth point cloud to compute the distance field ground truth numerically. We compute the distance field on a regular 3D grid of 5cm resolution for the RGB-D case and 10 cm for the LiDAR case. The RMSE error given the ground truth distance field is shown in Fig. LABEL:cow:distance and LABEL:newer:distance. Our distance field has the most accurate results. Voxblox produces the large distance RMSE for the RGB-D dataset, and VDB-EDT has a larger error on the Newer College dataset than others due to failure to clear the moving objects successfully.
To further examine the behaviour of VDB-GPDF, we compare 2D slices of the ground-truth and estimated distance fields in Fig. 6. We choose a horizontal slice $0.9$m above the ground, roughly cutting the scene objects in the middle. In addition, we evaluate the EDF with and without the G-GPDF enabled to show that G-GPDF produces more accurate distance inference than fusing the ESDF directly. Moreover, G-GPDF is a generative model that can generate distance and gradient inference even in places with no observations. As shown in Fig. LABEL:distance_c:edt,  LABEL:distance_c:vox and Fig. LABEL:distance_c:ours_direct_mode, VDB-EDT and Voxblox and VDB-GPDF without G-GPDF only compute the EDF within the sensor’s field of view, whereas VDB-GPDF with G-GPDF naturally predicts the EDF value at all points as can be seen in Fig. LABEL:distance_c:ours. For a quantitative evaluation, we illustrate the distance error with a colour map in Fig. LABEL:distance_c:error_edtLABEL:distance_c:error_voxLABEL:distance_c:error_ours_direct_mode and LABEL:distance_c:error_ours. Our VDB-GPDF produces lower and smoother values of distance errors.
Similarly, for the LiDAR dataset, we compare 2D slices of the ground-truth and estimated distance fields in Fig. 7. As shown in Fig. LABEL:distance_n:edtLABEL:distance_n:vox and LABEL:distance_n:ours, our VDB-GPDF computes the similar EDF as the ground truth in Fig. LABEL:distance_n:ground_truth. For a quantitative evaluation, we plot the same distance absolute error in Fig. LABEL:distance_c:error_edtLABEL:distance_c:error_vox and LABEL:distance_c:error_ours. Our VDB-GPDF generates lower and more consistent distance error values.
@@ VI Conclusion
UTS
University of Technology Sydney
RI
Robotics Institute
FEIT
Faculty of Engineering and Information Technology
CAS
Centre for Autonomous Systems
WiEIT
Women in Engineering and IT
1D
One-Dimensional
2D
Two-Dimensional
2.5D
Two-and-a-Half-Dimensional
3D
Three-Dimensional
GP
Gaussian Process
GPIS
Gaussian Process Implicit Surfaces
MAP
Maximum A Posteriori
MLE
Maximum Likelihood Estimation
ICP
Iterative Closest Point
MVE
Multi-View Environment
OG
Occupancy Grid
CHOMP
Covariant Hamiltonian Optimization for Motion Planning
FIESTA
Fast Incremental Euclidean DiSTAnce Fields
GPUs
Graphics Processing Units
KD
K-Dimensional
GPDF
Gaussian Process Distance Field
Log-GPIS
Log-Gaussian Process Implicit Surfaces
LiDAR
Light Detection And Ranging Sensor
SLAM
Simultaneous Localisation and Mapping
TDF
Truncated Distance Field
EDF
Euclidean Distance Field
SDF
Signed Distance Field
TSDF
Truncated Signed Distance Field
ESDF
Euclidean Signed Distance Field
GPIS-SDF
GPIS with signed distance function
RGB
Red-Green-Blue
RGB-D
Red-Green-Blue-Depth
RMSE
Root Mean Squared Error
CI
Conditional Independent
FP
Forward Propagation
BP
Backward Propagation
D-SKI
Structured Kernel Interpolation framework with Derivatives
SKI
Structured Kernel Interpolation method
D-SKI-CI-Fusion
Structured Kernel Interpolation with Derivatives and Conditional Independent Fusion
MVMs
Matrix-Vector Multiplications method
PDE
Partial Differential Equation
Log-GPIS-MOP
Log-Gaussian Process Implicit Surface for Mapping, Odometry and Planning
RANSAC
Random Sample Consensus
Dynamic-GPDF
Dynamic Gaussian Process Distance Fields

### v3

@@ IV-D Global Gaussian Process Signed Distance Field
After the fusion process, the leaf nodes in the global VDB grid are marked as active nodes. We perform the marching cubes for the voxels of active nodes and for computing the dense reconstruction with surface properties. The zero crossing points in current active nodes $\mathcal{P}_{\mathcal{S}\{i\}}$ then merge into $\mathcal{P}_{\mathcal{S}\{0,...,i\}}$. Zero crossing points in each node are used to train separate GPs. The combination of all GPs forms G-GPDF.
@@
where $\lambda$ is a large positive value to control the evenness of the smooth minimum. We use $100$ for $\lambda$ in this paper. ${\hat{d}}_{q}(\mathbf{x}_{*})$ is the distance inference at querying location $\mathbf{x}_{*}$ given the located GP node with the index of $q$. For the gradient, we directly average the inferred gradients from the $Q$ GP nodes.
@@ V Evaluation
To evaluate the proposed VDB-GPDF framework, we quantitatively compare the performance for a) fusion output, b) efficiency, c) reconstruction accuracy and d) distance accuracy. Our framework is implemented in C++ based on ROS1 with CPU only. All experiments were run on 12th Gen Intel® Core™ i5-1245U. We use frames 220-820 of the Cow and Lady dataset [1], which covers major information of the scene. For the real-world LiDAR dataset, we choose the quad of the Newer College dataset [24]. The estimated poses are calculated via the alignment of each frame to the ground truth map. The size of the main quad is roughly $56\times 40\times 20$ m, with approximately 2000 frames. We also include the larger simulated LiDAR dataset Mai City [25] sized $700\times 120\times 3$ m. All the above datasets have the ground truth map that allows us to perform quantitative evaluations for reconstruction and distance accuracy. In addition, we qualitatively validate our framework on the KITTI odometry 07 dataset [26] with dynamic objects in a large-scale map. Our strategy for length scale selection depends on the number of voxels in each leaf node, which in our case is fixed to $2^{3}\times 2^{3}\times 2^{3}$ voxels. A length scale that correlates two to four voxels is a good trade-off between smoothing and interpolation for our kernel function.
@@ V-B Efficiency
Figure 4: Efficiency statistical evaluation of the time per frame for the Cow and Lady (top), the Newer College (middle) and Mai City (bottom) datasets respectively.
We first evaluate the VDB-GPDF efficiency against the state-of-the-art mapping frameworks VDB-Fusion and VDB-EDT, both developed based on OpenVDB data structure, and Voxblox based on Hash maps. Fig. 4 shows the statistical evaluation of the computational time (mean and standard deviation) for multiple datasets with varying voxel sizes. Note that the Newer College quad has pedestrians walking by as dynamic objects. All frameworks have free-space carving methods enabled to update the moving objects and noise in the scene. For the sake of simplicity, we compute the time for all the integration processes, including fusion and ESDF. As for our framework, we include L-GPDF, incremental fusion and G-GPDF. For the Cow and Lady, Voxblox is overall more efficient than others while our VDB-GPDF outperforms Voxblox in high resolution. For LiDAR cases, our VDB-GPDF is more efficient than other frameworks with varying resolutions due to our voxelisation method and fewer testing points in the frustum area.
Figure 5: Qualitative reconstruction comparison of a) VDBFusion and b) our proposed method shows that VDB-GPDF produces a more complete and natural reconstruction, especially when the sensor is in parallel with the observed surface. Our mesh covers on the top of the mattress, and the colour is properly fused for visualisation.
Figure 6: Quantitative comparisons of a), c) and e) reconstruction accuracy in Chamfer distance, and b), d) and f) the distance field accuracy in RMSE for the Cow and Lady (top), the Newer College (middle) and Mai City (bottom) datasets respectively.
@@ V-C Reconstruction Accuracy
For a fair comparison, we process every frame in each dataset for each framework. VDBFusion and VDB-GPDF both produce similar meshes of the scene shown in Fig. LABEL:cow_mesh_vdbfusion and Fig. LABEL:cow_mesh_ours, respectively. VDB-GPDF produces a more complete and natural reconstruction, especially when the sensor is parallel to the observed surface. It can be seen that there is a gap at the top of the mattresses on the left-hand side in the VDBFusion result, whereas VDB-GPDF does not have the same gap. In addition, our colour is nicely fused with a sample of the RGB camera data for qualitative comparison.
We compare the maps on both datasets quantitatively using the Chamfer distance metrics [27] against VDB-Fusion and VDB-EDT. As demonstrated in Fig. LABEL:cow:chamfer on the Cow and Lady, LABEL:newer:chamfer Newer college and LABEL:mai:chamfer Mai City, our map quality outperforms VDB-EDT in different resolutions. Due to our voxelisation method being applied, we have comparable results to VDBFusion in reconstruction accuracy. In Fig. 1 on Newer College, we qualitatively show our mesh coloured by fused LiDAR intensity and the distance field output. Fig. LABEL:teaser1 shows the incrementally built dense reconstruction. We zoomed in Fig. LABEL:teaser2 to show the textures on the ceiling inside the corridor and Fig. LABEL:teaser3 the stairs and windows at the corner of the quad. Note that there exists a trade-off between accuracy and efficiency. For instance, efficiency decreases, and accuracy increases when more training points are added in L-GPDF and denser testing points are used in the frustum. In the results presented here, we achieve comparable reconstruction and distance accuracy with better time consumption.
Figure 7: Estimated EDF of cow and lady in e) is accurate and closely resembles the ground truth in a). Our VDB-GPDF with G-GPDF in i) produces lower and smoother errors than others.
Figure 8: Estimated EDF of the newer college in d) is accurate and closely resembles the ground truth in a). b) fails to clean the areas where moving objects are present. c) does not produce enough smoothness in the field. Similarly, Our VDB-GPDF in h) produces lower values of distance errors.
@@ V-D Distance Accuracy
We evaluate the proposed GP distance field with VDB structure against Voxblox and VDB-EDT since the VDBFusion framework has no EDF. We use the ground truth point cloud to compute the distance field ground truth numerically. We compute the distance field on a regular 3D grid of 5cm resolution for the RGB-D case and 10 cm for the LiDAR case. The RMSE error given the ground truth distance field is shown in Fig. LABEL:cow:distanceLABEL:newer:distance and  LABEL:mai:distance. Our distance field has the most accurate results. Voxblox produces the large distance RMSE for the Cow and Lady and Mai City datasets, and VDB-EDT has a larger error on the Newer College dataset than others due to failure to clear the moving objects successfully.
To further examine the behaviour of VDB-GPDF, we compare 2D slices of the ground-truth and estimated distance fields for the Cow and Lady in Fig. 7. We choose a horizontal slice $0.9$m above the ground, roughly cutting the scene objects in the middle. In addition, we evaluate the EDF with and without the G-GPDF enabled to show that G-GPDF produces more accurate distance inference than fusing the ESDF directly. Moreover, G-GPDF is a generative model that can generate distance and gradient inference even in places with no observations. As shown in Fig. LABEL:distance_c:edt,  LABEL:distance_c:vox and Fig. LABEL:distance_c:ours_direct_mode, VDB-EDT and Voxblox and VDB-GPDF without G-GPDF only compute the EDF within the sensor’s field of view, whereas VDB-GPDF with G-GPDF naturally predicts the EDF value at all points as can be seen in Fig. LABEL:distance_c:ours. For a quantitative evaluation, we illustrate the distance error with a colour map in Fig. LABEL:distance_c:error_edtLABEL:distance_c:error_voxLABEL:distance_c:error_ours_direct_mode and LABEL:distance_c:error_ours. Our VDB-GPDF produces lower and smoother values of distance errors. Similarly, as shown in Fig. LABEL:distance_n:edtLABEL:distance_n:vox and LABEL:distance_n:ours, our VDB-GPDF computes the similar EDF as the ground truth in Fig. LABEL:distance_n:ground_truth. For a quantitative evaluation, we plot the same distance absolute error in Fig. LABEL:distance_n:edt_errorLABEL:distance_n:vox_error and LABEL:distance_n:ours_error. Our VDB-GPDF generates lower and more consistent distance error values.
V-E Dynamic Performance
We evaluate our framework with and without the free space carving method on the large-scale KITTI dataset that contains multiple dynamic objects. We compare against the ground truth static map generated by [28] and colour the vertices of the reconstructed mesh with error. The black sections in Fig. LABEL:kitti_1 indicate the map has accumulated defects due to the moving cars and pedestrians. The dynamic objects have been updated in the scene as shown in Fig. LABEL:kitti_2.
Figure 9: Kitti dataset map coloured by error (red $\downarrow$ and green $\uparrow$) a) without space carving and b) with space carving.